if R and Hare the horizontal range and maximum hieght attained by a projectile, then its speed for projection is?

Grade:12th pass

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The Range and vertical height would be given by the relation:
R=\frac{u^2\sin2\theta}{g}
H=\frac{u^2\sin^2\theta}{2g}
Hence,
\frac{R}{H}=\frac{u^2\sin2\theta}{g}\frac{2g}{u^2\sin^2\theta}=4\tan\theta
Using this, we can find the speed of projection.
Regards
shrilekha sharma
9 years ago
sir but the answer given in the key is under root 2gH +R2g/8H sir please explain in detail?